OpenStudy (anonymous):

Two of your friends, Matt and Karen, both run to you to settle a dispute. They were working on a math problem, and got different answers. Wisely, you decide to look at their work to see if you can spot the source of confusion. Matt 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(4) + 30 ÷ 5 6 – 16 + 30 ÷ 5 −10 + 30 ÷ 5 20 ÷ 5 4 Karen 6 – 4(3 – 5)2 + 30 ÷ 5 6 – 4(–2)2 + 30 ÷ 5 6 – 4(−4) + 30 ÷ 5 6 + 16 + 30 ÷ 5 6 + 16 + 6 22 + 6 28 Explain to Matt and Karen who, if either, is correct, and identify error

OpenStudy (tkhunny):

(-2)^2 = +4

OpenStudy (lanhikari22):

dude, please when dealing with exponents don't do things like 6 – 4(3 – 5)2 + 30 ÷ 5 either do 6 – 4(3 – 5)^2 + 30 ÷ 5 or 6 – 4(3 – 5)² + 30 ÷ 5

OpenStudy (mathstudent55):

They both made mistakes, so both of their answers are wrong.

OpenStudy (mathstudent55):

Matt \(6 – 4(3 – 5)^2 + 30 ÷ 5 \) \(= 6 – 4(–2)^2 + 30 ÷ 5 \) \(=6 – 4(4) + 30 ÷ 5 \) \(=6 – 16 + 30 ÷ 5 \) \(=−10 + \color{red}{30 ÷ 5} \) \(=-10 + \color{red}{6} \) \(=-4 \) Karen \(6 – 4(3 – 5)^2 + 30 ÷ 5 \) \(=6 – 4\color{red}{(–2)^2} + 30 ÷ 5 \) \(=6 – 4\color{red}{(4)} + 30 ÷ 5 \) \(=6 - 16 + 30 ÷ 5 \) \(=6 - 16 + 6 \) \(=-10 + 6 \) \(=-4\)